Find the area of the triangle formed with the co-ordinate axes and the tangent drawn at the point (3, -4) on the circle x2 + y2 =25 is

Grade:11

2 Answers

Khimraj
5 years ago
x2 + y2 = 25
differentiating wrt x
2x + 2y.dy/dx = 0
dy/dx = -x/y
slope m = -3/-4 = ¾
so equation of tangent
y+4 = (3/4)*(x-3)
it cuts x -axis where y=0
so x = 16/3 + 3 = 25/3
it cuts y -axis where x=0
y = -25/4
so area of truangle = ½*(25/3)(25/4) = 625/24
Hope it clears..................
AKSHANSH PAREEK
5 years ago
draw perpendicular on tangent this will be height of triangle=5(radius)  
find length of tangent this will be bae of trianle 
now use ½ base* height uwiill get area of triangle